# Variable Selection in Convex Piecewise Linear Regression

### v1

@@ 2.1 Sparse Gradient Descent Algorithm
where ${\left\lVert\cdot\right\rVert}_{0}$ counts the number of nonzero entries and $[\bm{\varphi}]_{1:d}$ denotes the sub-vector of $\bm{\varphi}\in\mathbb{R}^{d+1}$ with the first $d$ entries.
@@
${}~{}\bm{\theta}_{j}^{t+1}\leftarrow\bm{\Psi}_{s}(\bm{\alpha}_{j}^{t+1})$
@@
                   ${}~{}\bm{\alpha}_{j}^{t+1}\leftarrow\bm{\theta}_{j}^{t}$
@@
Sp-GD is a variant of the projected gradient descent algorithm to pursue the above estimators, where the gradient is substituted by the generalized gradient [Hiriart-Urruty, 1979] and the step size varies across blocks adaptively with the iterates. We introduce a geometric object to describe the Sp-GD algorithm. Let
[…]
@@
The max-affine function in (6) is a special instance of tropical polynomials in the max-plus algebra [Maragos et al., 2021]. In this perspective, the sets $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\mathcal{V}(\bm{\theta})$ are called respectively as tropical open cells and tropical zero-set.
Algorithm 1 presents a pseudo code for the SpGD algorithm. Each iteration of the algorithm starts by updating the empirical probability $\pi_{j}^{t}$ that the covariates belong to then open set $\mathcal{C}_{j}(\bm{\theta}^{t})$ determined by the previous iterate $\bm{\theta}^{t}$ for all $j\in[k]$. Note that the evaluation of the indicator function in (9) can be saved and reused in the subsequent step in (10). If $\pi_{j}^{t}$ is non-zero, then we update the $j$th block by generalized gradient descent in (10) followed by the orthogonal projection to $\Gamma_{s}$ given by
@@
Otherwise, the $j$th block remains the same as the previous iterate. This update rule applies recursively until the algorithm converges by satisfying $\lVert\bm{\theta}^{t+1}-\bm{\theta}^{t}\rVert_{2}/\lVert\bm{\theta}^{t}\rVert_{2}$ less than a given threshold. Note that the step size for the gradient descent is adaptively evaluated as the reciprocal of $\pi_{j}^{t}$ which is determined solely by the previous iterate $\bm{\theta}^{t}$. In other words, this algorithm does not require tuning for the step size. In fact, the step size in (10) is always larger than $1$, which is quite different from the typical choices of step size for gradient descent (e.g. a small constant or a diminishing sequence). The next section presents theoretical guarantees of Sp-GD with this proposed step size. Furthermore, as shown in Section 4.1, this step size rule also makes Sp-GD empirically converge fast to the desired estimate.
@@ 2.2 Theoretical Analysis of Sp-GD
$\mathcal{N}(\bm{\theta}^{\star},\kappa\rho):=\Bigg{\{}\bm{\theta}\in\mathbb{R}^{k(d+1)}\,:\,{\left\lVert\bm{\theta}-\bm{\theta}^{\star}\right\rVert}_{2}\leq\kappa\rho\Bigg{\}}$ (15)
@@
Theorem 2.3 implies local linear convergence of Sp-GD in the noiseless case when the algorithm is properly initialized for the sub-Gaussian covariate model. The sample complexity scales linearly with $s$ significantly improving analogous results without variable selection [Ghosh et al., 2021, Kim and Lee, 2024]. Importantly, Sp-GD does not inflate the degree of dependence on the model order $k$ and dataset imbalance parameter $\pi_{\min}$ and maintains the same order $k^{4}\pi_{\min}^{-12}$ as plain GD and SGD [Kim and Lee, 2024]. Therefore, Sp-GD outperforms these algorithms even in the absence of sparsity in the max-affine model.
@@ 3.1 Initialization Algorithms
To bypass this issue, we compute an initial estimate in the basin of attraction of the ground truth so that the subsequent Sp-GD converges to the desired estimate. One may apply the initialization scheme for max-affine regression by Ghosh et al. [2021] while ignoring the sparsity constraint. However, their theoretical guarantees provide the desired accuracy when the sufficient number of observations scales at least linearly in the total number of variables $d$. Therefore, using this initialization scheme with Sp-GD yields a requirement on the sample complexity that is no longer dominated by the number of active variables $s$. To retain the gain due to Theorem 2.3, we propose an initialization scheme modified from the spectral initialization for max-affine regression [Ghosh et al., 2021] that provides an error rate depending on $s$ instead of $d$. To fulfill this objective, the initialization scheme additionally requires that the $s$-sparse weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ are jointly supported on a set with cardinality at most $s$. In other words, the ground-truth parameter vector $\bm{\theta}^{\star}=[[\bm{a}_{1}^{\star};b_{1}^{\star}];\ldots;[\bm{a}_{k}^{\star};b_{k}^{\star}]]$ belongs to the set defined by
@@
$\bm{\widehat{m}}_{1}\leftarrow\sum_{i=1}^{n}y_{i}\bm{x}_{i},$ $\quad\bm{\widehat{M}}_{2}\leftarrow\sum_{i=1}^{n}y_{i}\left(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d}\right),$ $\quad\bm{\widehat{M}}\leftarrow\bm{\widehat{m}}_{1}\bm{\widehat{m}}_{1}^{\scriptscriptstyle{\textup{{T}}}}+\widehat{\bm{M}}_{2}$
$\bm{\widehat{P}}\leftarrow\displaystyle\mathop{\mathrm{argmax}}_{\widetilde{\bm{P}}\in\mathcal{F}^{k}_{d}}~{}\mathrm{tr}\left(\bm{\widehat{M}}^{\scriptscriptstyle{\textup{{T}}}}\widetilde{\bm{P}}\right)-\lambda\sum\limits_{i,j}\left|[\widetilde{\bm{P}}]_{i,j}\right|$triangleright where $\mathcal{F}^{k}_{d}$ is defined in (21)
$\mathcal{S}\leftarrow\left\{\text{indices of $s$-largest diagonal entries of $\widehat{\bm{P}}$}\right\},\quad\bm{\widehat{V}}\leftarrow\displaystyle\mathop{\operatorname*{argmin}}_{\widetilde{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}\widetilde{\bm{V}}=\bm{I}_{k}}{\left\lVert[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}-\widetilde{\bm{V}}\widetilde{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}\right\rVert}_{\mathrm{F}}$
@@
First, the subspace estimation algorithm is presented in Algorithm 2. This algorithm modifies upon the moment method for parameter estimation originally developed for mixtures of linear models [Chaganty and Liang, 2013, Zhang et al., 2014, Yi et al., 2016, Sedghi et al., 2016]. In Algorithm 2, $\widehat{\bm{m}}_{1}$ and $\widehat{\bm{M}}_{2}$ are respectively the first and second central moments weighted by the target values. In the non-sparse case, applying PCA to $\widehat{\bm{M}}$ and taking the first $k$-dominant eigenvectors yields a basis estimate of the subspace spanned by the ground-truth weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$. In our case, the joint sparsity of the dominant eigenvectors can be utilized to obtain a more accurate estimate of the subspace via sparse PCA (sPCA) [Zou et al., 2006]. Several convex relaxations were proposed to solve the sPCA problem [d’Aspremont et al., 2008, Amini and Wainwright, 2009, Zhang et al., 2012, Vu et al., 2013, Dey et al., 2018, Li and Xie, 2020]. Algorithm 2 uses the convex relaxation of sPCA by [Vu et al., 2013] that computes the $k$-dominant eigenvectors under the joint sparsity constraint. They formulated the estimation of sparse principal components as the following semidefinite program:
@@
Indeed, the Fantope $\mathcal{F}^{k}_{d}$ is obtained as the convex hull of all rank $k$-projection matrices [Overton and Womersley, 1992]. On one hand, the first term in the maximization in (20) measures the similarity between an element in the Fantope and the empirical moment matrix $\widehat{\bm{M}}$. On the other hand, the second term in (20) encourages sparsity by penalizing with the $\ell_{1}$ norm of all entries of the matrix. This relaxed version of sPCA is applied in the second step of Algorithm 2. The solution is calculated using the Alternating Direction Method of Multipliers (ADMM) algorithm presented in [Vu et al., 2013, Algorithm 1]. Since the result $\widehat{\bm{P}}$ is not necessarily a valid rank-$k$ projection matrix, we need post-processing in the following two steps. First, the algorithm recovers the support as $\mathcal{S}$ by thresholding the diagonal entries of $\widehat{\bm{P}}$. The final step in Algorithm 2 is to find the optimal projection of $[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}$ onto the set of all rank-$k$ projection matrices, which can be obtained by the $k$-dominant eigenvectors of $[\widehat{\bm{P}}]_{\mathcal{S},\mathcal{S}}$. Then $\widehat{\bm{V}}\in\mathbb{R}^{s\times k}$ will denote the Cholesky factor of the estimated rank-$k$ projection matrix.
@@
$\mathcal{N}\leftarrow r$-covering of $B_{2}^{k+1}$
($\left(\{[\bm{w}_{j}^{\sharp};b_{j}^{\sharp}]\}_{j=1}^{k},c^{\sharp}\right)\leftarrow\mathop{\mathrm{argmin}}_{\{[\bm{w}_{j};b_{j}]\}_{j=1}^{k}\in\mathcal{N},c\geq 0}\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{i}-c\cdot\underset{j\in[k]}{\max}\langle[\bm{x}_{i}]_{\mathcal{S}}^{\mathrm{T}}\widehat{\bm{V}}w_{j}+b_{j}\rangle\right)^{2}$
@@
       $[\bm{\theta}_{j}^{0}]_{\mathcal{S}}\leftarrow c^{\sharp}[\widehat{\bm{V}}\bm{w}_{j}^{\sharp};b_{j}^{\sharp}],\quad[\bm{\theta}_{j}^{0}]_{[d]\setminus\mathcal{S}}\leftarrow\bm{0}\in\mathbb{R}^{d-s}$
@@
Algorithm 3 Discrete Search over Estimated Subspace [Ghosh et al., 2021]
Algorithm 2 identifies the joint support of $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and estimates the subspace spanned by these vectors. This information is not sufficient to estimate the individual parameter vectors $\{[\bm{a}_{j}^{\star},b_{j}^{\star}]\}_{j=1}^{k}$. To approximate the parameter vectors up to a global scaling, Ghosh et al. [2021] proposed a discrete search over subsets of an $r$-covering $\mathcal{N}$ of the unit ball that satisfies $\min_{\bm{x}\in\mathcal{N}}\|\bm{x}-\bm{u}\|\leq r$ for all $\bm{u}\in B_{2}^{k+1}$ and $B_{2}^{k+1}\subset\bigcup_{\bm{w}\in\mathcal{N}}\left(\bm{w}+rB_{2}^{k+1}\right)$. The cost of constructing an $r$-covering and searching over it grows exponentially in the dimension. Due to the dimensionality reduction by Algorithm 2, the search dimension is the model order $k$, which is often much smaller than the ambient dimension $d$. If this is the case, then the discrete search is computationally feasible. This method also applies to the initialization of sparse max-affine regression. To make the manuscript self-contained, we summarize the discrete search algorithm by Ghosh et al. [2021] in Algorithm 3 using our notation.

### v2

@@ 2.1 Sparse Gradient Descent Algorithm
where ${\left\lVert\cdot\right\rVert}_{0}$ counts the number of nonzero entries and $\bm{\varphi}]_{1:d}$ denotes the sub-vector of $\bm{\varphi}\in\mathbb{R}^{d+1}$ with the last entry omitted.
@@
$\bm{\theta}_{j}^{t+1}\leftarrow\bm{\Psi}_{s}(\bm{\alpha}_{j}^{t+1})$
@@
       $\bm{\alpha}_{j}^{t+1}\leftarrow\bm{\theta}_{j}^{t}$
@@
Sp-GD is a variant of the projected gradient descent algorithm to pursue the above estimators. In Sp-GD the gradient is substituted by the generalized gradient (Hiriart-Urruty, 1979) and the step size varies across blocks adaptively with the iterates. We introduce a geometric object to describe the Sp-GD algorithm. Let
$\mathcal{C}_{j}([\bm{\theta}_{1};\dots;\bm{\theta}_{k}])\triangleq\left\{\bm{x}\in\bm{b}{R}^{d}~:~\langle[\bm{x};\ 1],\bm{\theta}_{j}\rangle>\langle[\bm{x};\ 1],\bm{\theta}_{l}\rangle,~\forall l\neq j\right\}$ (14)
denote an open set in $\mathbb{R}^{d}$ where the $j$th linear model $\bm{x}\mapsto\langle[\bm{x};\ 1],\bm{\theta}_{j}\rangle$ achieves the unique maximum for $j\in[k]$. The ties in the maximum occur on the set
$\mathcal{V}([\bm{\theta}_{1};\dots;\bm{\theta}_{k}])\triangleq\bigcup_{l\neq j}\left\{\bm{x}\in\bm{b}{R}^{d}~:~\langle[\bm{x};\ 1],\bm{\theta}_{j}\rangle=\langle[\bm{x};\ 1],\bm{\theta}_{l}\rangle\right\}.$
@@
The max-affine function in (8) is a special instance of tropical polynomials in the max-plus algebra (Maragos et al., 2021). In this perspective, the sets $\{\mathcal{C}_{j}(\bm{\theta})\}_{j=1}^{k}$ and $\mathcal{V}(\bm{\theta})$ are called respectively as tropical open cells and tropical zero set.
Algorithm 1 presents a pseudo code for the Sp-GD algorithm. Each iteration of the algorithm starts by updating the empirical probability $\pi_{j}^{t}$ that the covariates belong to the open set $\mathcal{C}_{j}(\bm{\theta}^{t})$ determined by the previous iterate $\bm{\theta}^{t}$ for all $j\in[k]$. Note that the evaluation of the indicator function in (11) can be saved and reused in the subsequent step in (12). If $\pi_{j}^{t}$ is non-zero, then we update the $j$th block by generalized gradient descent in (12) followed by the orthogonal projection to $\Gamma_{s}$ given by
@@
Otherwise, the $j$th block remains unchanged from the previous iterate. This update rule applies recursively until the algorithm converges by satisfying $\lVert\bm{\theta}^{t+1}-\bm{\theta}^{t}\rVert_{2}/\lVert\bm{\theta}^{t}\rVert_{2}$ less than a given threshold. Note that the step size for the gradient descent is adaptively evaluated as the reciprocal of $\pi_{j}^{t}$ which is determined solely by the previous iterate $\bm{\theta}^{t}$. In other words, this algorithm does not require tuning for the step size. In fact, the step size in (12) is always larger than $1$, which is quite different from the typical choices of step size for gradient descent (e.g. a small constant or a diminishing sequence). The next section presents theoretical guarantees of Sp-GD with this proposed step size. Furthermore, as shown in Section 5.1, this step size rule also makes Sp-GD empirically converge fast to the desired estimate.
@@ 2.2 Theoretical Analysis of Sp-GD
$\mathcal{N}(\bm{\theta}^{\star},\kappa\rho)=\Bigg\{\bm{\theta}\in\mathbb{R}^{k(d+1)}\,:\,{\left\lVert\bm{\theta}-\bm{\theta}^{\star}\right\rVert}_{2}\leq\kappa\rho\Bigg\}$ (17)
@@
Theorem 2.3 implies local linear convergence of Sp-GD in the noiseless case when the algorithm is properly initialized for the sub-Gaussian covariate model. The sample complexity scales linearly with $s$ significantly improving analogous results without variable selection (Ghosh et al., 2021; Kim and Lee, 2024). Importantly, Sp-GD does not inflate the degree of dependence on the model order $k$ and dataset imbalance parameter $\pi_{\min}$ and maintains the same order $k^{4}\pi_{\min}^{-12}$ as plain GD and SGD (Kim and Lee, 2024). Therefore, Sp-GD outperforms these algorithms even in the absence of sparsity in the max-affine model.
@@ 3.1 Initialization Algorithms
To bypass this issue, we compute an initial estimate in the basin of attraction of the ground truth so that the subsequent Sp-GD converges to the desired estimate. One may apply the initialization scheme for max-affine regression by (Ghosh et al., 2021) while ignoring the sparsity constraint. However, their theoretical guarantees provide the desired accuracy when the sufficient number of observations scales at least linearly in the total number of variables $d$. Therefore, using this initialization scheme with Sp-GD yields a requirement on the sample complexity that is no longer dominated by the number of active variables $s$. To retain the gain due to Theorem 2.3, we propose an initialization scheme modified from the spectral initialization for max-affine regression (Ghosh et al., 2021) that provides an error rate depending on $s$ instead of $d$. To fulfill this objective, the initialization scheme additionally requires that the $s$-sparse weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ are jointly supported on a set with cardinality at most $s$. In other words, the ground-truth parameter vector $\bm{\theta}^{\star}=[[\bm{a}_{1}^{\star};b_{1}^{\star}];\ldots;[\bm{a}_{k}^{\star};b_{k}^{\star}]]$ belongs to the set defined by
@@
0.05in $\bm{\widehat{m}}_{1}\leftarrow\frac{1}{n}\sum_{i=1}^{n}y_{i}\bm{x}_{i},$ $\quad\bm{\widehat{M}}_{2}\leftarrow\frac{1}{n}\sum_{i=1}^{n}y_{i}\left(\bm{x}_{i}\bm{x}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d}\right),$ $\quad\bm{\widehat{M}}\leftarrow\bm{\widehat{m}}_{1}\bm{\widehat{m}}_{1}^{\scriptscriptstyle{\textup{{T}}}}+\widehat{\bm{M}}_{2}$
0.05in $\bm{\widehat{P}}\leftarrow\displaystyle\mathop{\mathrm{argmax}}_{\widetilde{\bm{P}}\in\mathcal{F}^{k}_{d}}~\mathrm{tr}\left(\bm{\widehat{M}}^{\scriptscriptstyle{\textup{{T}}}}\widetilde{\bm{P}}\right)-\lambda\sum\limits_{i,j}\left|[\widetilde{\bm{P}}]_{i,j}\right|$ $\triangleright$ where $\mathcal{F}^{k}_{d}$ is defined in (23)
0.05in $\mathcal{S}\leftarrow\left\{\text{indices of $s$-largest diagonal entries of $\widehat{\bm{P}}$}\right\},\quad\bm{\widehat{V}}\leftarrow\displaystyle\mathop{\operatorname*{argmin}}_{\widetilde{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}\widetilde{\bm{V}}=\bm{I}_{k}}{\left\lVert[\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}-\widetilde{\bm{V}}\widetilde{\bm{V}}^{\scriptscriptstyle{\textup{{T}}}}\right\rVert}_{\mathrm{F}}$
@@
First, the subspace estimation algorithm is presented in Algorithm 2. This algorithm modifies upon the moment method for parameter estimation originally developed for mixtures of linear models (Chaganty and Liang, 2013; Zhang et al., 2014; Yi et al., 2016; Sedghi et al., 2016). In Algorithm 2, $\widehat{\bm{m}}_{1}$ and $\widehat{\bm{M}}_{2}$ are respectively the first and second central moments weighted by the target values. In the non-sparse case, applying PCA to $\widehat{\bm{M}}$ and taking the first $k$-dominant eigenvectors yields a basis estimate of the subspace spanned by the ground-truth weight vectors $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$. In our case, the joint sparsity of the dominant eigenvectors can be utilized to obtain a more accurate estimate of the subspace via sparse PCA (sPCA) (Zou et al., 2006). Several convex relaxations were proposed to solve the sPCA problem (d’Aspremont et al., 2008; Amini and Wainwright, 2009; Zhang et al., 2012; Vu et al., 2013; Dey et al., 2018; Li and Xie, 2020). Algorithm 2 uses the convex relaxation of sPCA by (Vu et al., 2013) that computes the $k$-dominant eigenvectors under the joint sparsity constraint. They formulated the estimation of sparse principal components as the following semidefinite program:
@@
Indeed, the Fantope $\mathcal{F}^{k}_{d}$ is obtained as the convex hull of all rank $k$-projection matrices (Overton and Womersley, 1992). On one hand, the first term in the maximization in (22) measures the similarity between an element in the Fantope and the empirical moment matrix $\widehat{\bm{M}}$. On the other hand, the second term in (22) encourages sparsity by penalizing with the $\ell_{1}$ norm of all entries of the matrix. This relaxed version of sPCA is applied in the second step of Algorithm 2. The solution to this is calculated using the Alternating Direction Method of Multipliers (ADMM) algorithm presented in (Vu et al., 2013, Algorithm 1). Since the result $\widehat{\bm{P}}$ is not necessarily a valid rank-$k$ projection matrix, we need post-processing in the following two steps. First, the algorithm recovers the support as $\mathcal{S}$ by thresholding the diagonal entries of $\widehat{\bm{P}}$. The final step in Algorithm 2 is to find the optimal projection of $\bm{\widehat{P}}]_{\mathcal{S},\mathcal{S}}$ onto the set of all rank-$k$ projection matrices, which can be obtained by the $k$-dominant eigenvectors of $\widehat{\bm{P}}]_{\mathcal{S},\mathcal{S}}$. Then $\widehat{\bm{V}}\in\mathbb{R}^{s\times k}$ will denote the Cholesky factor of the estimated rank-$k$ projection matrix.
@@
0.05in $\mathcal{N}\leftarrow r$-covering of $B_{2}^{k+1}$
0.05in $\left(\{[\bm{w}_{j}^{\sharp};b_{j}^{\sharp}]\}_{j=1}^{k},c^{\sharp}\right)\leftarrow\mathop{\mathrm{argmin}}_{\{[\bm{w}_{j};b_{j}]\}_{j=1}^{k}\in\mathcal{N},c\geq 0}\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{i}-c\cdot\underset{j\in[k]}{\max}\langle[\bm{x}_{i}]_{\mathcal{S}}^{\mathrm{T}}\widehat{\bm{V}}w_{j}+b_{j}\rangle\right)^{2}$
@@
 $\bm{\theta}_{j}^{0}]_{\mathcal{S}}\leftarrow c^{\sharp}[\widehat{\bm{V}}\bm{w}_{j}^{\sharp};b_{j}^{\sharp}],\quad[\bm{\theta}_{j}^{0}]_{[d]\setminus\mathcal{S}}\leftarrow\bm{0}\in\mathbb{R}^{d-s}$
@@
0.1in
Algorithm 3 Discrete Search over Estimated Subspace (Ghosh et al., 2021)
Algorithm 2 identifies the joint support of $\{\bm{a}_{j}^{\star}\}_{j=1}^{k}$ and estimates the subspace spanned by these vectors. This information is not sufficient to estimate the individual parameter vectors $\{[\bm{a}_{j}^{\star},b_{j}^{\star}]\}_{j=1}^{k}$. To approximate the parameter vectors up to a global scaling, the authors in (Ghosh et al., 2021) proposed a discrete search over subsets of an $r$-covering $\mathcal{N}$ of the unit ball that satisfies $\min_{\bm{x}\in\mathcal{N}}\|\bm{x}-\bm{u}\|\leq r$ for all $\bm{u}\in B_{2}^{k+1}$ and $B_{2}^{k+1}\subset\bigcup_{\bm{w}\in\mathcal{N}}\left(\bm{w}+rB_{2}^{k+1}\right)$. The cost of constructing an $r$-covering and searching over it grows exponentially in the dimension. Due to the dimensionality reduction by Algorithm 2, the search dimension is the model order $k$, which is often much smaller than the ambient dimension $d$. If this is the case, then the discrete search is computationally feasible. This method also applies to the initialization of sparse max-affine regression. To make the manuscript self-contained, we summarize the discrete search algorithm by (Ghosh et al., 2021) in Algorithm 3 using our notation.